# Advocating for the Silent: Enhancing Federated Generalization for Non-Participating Clients

### v4

@@ I Introduction
Federated Learning (FL) offers a collaborative paradigm to train a shared global model across distributed clients, ensuring data privacy by eliminating the need for direct data transfers [1, 2, 3]. Therefore, FL provides a secure architecture to bolster confidentiality and effectively manage diverse sensitive data [4], ranging from financial to healthcare information. Consequently, FL applications span various domains, encompassing finance [5], healthcare [4], recommendation systems [6], Internet-of-Things (IoT) [7], and more. However, the inherent heterogeneity among clients—often due to their distinct operational environments—generates Non-Independent and Identically Distributed (Non-IID) scenario [8, 9, 10]. This distinctness complicates the assessment of FL’s generalization capabilities, setting it apart from traditional centralized learning [11, 12].
@@
Motivation. Compared with traditional FL deploying a global model for actively participating clients and assuming an alignment between training and testing data distributions, our investigation spotlights the global model’s out-of-distribution generalization capabilities when trained by participating clients. In essence, we seek to cultivate a model via FL across active clients that reliably serves even the passive ones. Specifically, a novel contribution of our paper is the introduction of the self-information weighted expected risk – a metric to gauge model generalization. Our hypothesis is grounded in the conception that a model exhibiting proficiency with low-probability examples from training distributions might demonstrate adaptability to unfamiliar testing distributions. Such examples are anticipated to hold a greater significance in these unseen datasets. Operationalizing this concept within FL, our framework leverages the self-information of examples to craft a generalization boundary. Delving into the resulting information entropy-aware and distribution discrepancy-aware generalization disparities, a revelation emerges: data sources tend to manifest informational redundancy. This implies that certain data sources, characterized by diminished informational weight, can be seamlessly supplanted by others. Essentially, only a fraction of the client base holds substantial sway over FL’s generalization. This redundancy is especially pronounced in the IoT scenario, where a myriad of edge devices—often operating in overlapping zones—partake in FL. Consider drones, serving as FL clients, and amassing spatial data for model training. Drones constrained by limited flight spans might be rendered redundant due to their shared operational territories with other drones [19]. Informed by these insights, our paper further proposes strategies to enhance the generalization capabilities of FL, enabling trained models to better serve non-participating clients.
@@ II Related Work
Besides, compared with Federated Domain Generalization (FedDG) [27, 28] dealing with the challenge of domain distribution shift under FL settings, our focus is solely on a specific collection of all the data sources in FL. Our objective is to train a global model from seen distributions of participating clients and generalize it to other unseen distributions of non-participating clients. Additionally, it is commonly assumed that each domain contains equal information and cannot represent each other in FedDG. However, we consider the presence of information redundancy among data sources and only a subset of clients contributes to the generalization of FL. Most DG studies concentrate on learning an invariant representation across different domains [29, 30, 31]. Conversely, our focus is to design proper weighting aggregation and client selection methods to mitigate the generalization error in FL.
@@ III Theoretical Framework
Similar to previous studies [9, 11], we model each data source as a random variable with its corresponding distribution. In order to better evaluate the generalization performance of FL to ensure trained models perform well for non-participating clients, we focus on the information-theoretic generalization gap in FL. To conserve space, we have provided proofs for all our theorems in the appendix.
@@ III-A Preliminaries
The local training set $S_{i}=\{s_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in\mathcal{I}_{p}$ stored on participating client $i$ is made of $n_{i}$ i.i.d. realizations from the local data source $Z_{i}\sim P_{Z_{i}}$. Referring to [11], the objective of federated generalization is to train a global model on $\{S_{i}\}_{i\in\mathcal{I}_{p}}$, such that all the possible clients in $\mathcal{I}$ will be provided satisfactory service by this global model trained by participating clients. Let $\mathcal{H}$ be a hypothesis class on $\mathcal{X}$. The loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ is a non-negative function and we assume that $\ell$ is bounded by $b$. For simplicity, we denote $\ell(h(x),y)$ as $\ell(h,z)$ in the following.
@@
The rationale behind the formulation of the risk outlined in (1) can be attributed to the fact that target distributions are unknown in the training stage under the OOD setting. The proposed loss $\ell(h,z_{i})\log(\frac{1}{P_{Z_{i}}(z_{i})})$, is indicative of the requisite focus on outcomes $z_{i}\in\mathcal{Z}$ with lower probabilities in source distributions since they may have higher probabilities in unknown target distributions. Additionally, it is reasonable to assume that unknown data sources have maximum entropy distributions, signifying greater uncertainty, exemplified by a uniform distribution of discrete labels. For example, IoT devices may only collect data in specific areas, while the global model trained by these devices is expected to provide spatial-related services for all devices across the entire area. Hence, the self-information of each outcome is indispensable for measuring the expected risk in such a situation.
@@ III-B Federated Generalization
Theorem 1 asserts that increasing the weighted sum $\sum_{i\in\mathcal{I}_{p}}\alpha_{i}H(Z_{i})$ will reduce this generalization bound. This suggests that models trained by FL will exhibit enhanced performance on unknown data sources when a greater weighting factor is assigned to the data source with richer information. In addition, model complexity $VC(\mathcal{G})$ and sample complexity $\sum_{i=1}^{|\mathcal{I}_{p}|}n_{i}$ also affect the generalization capacity of FL.
We then consider assigning the identical weighting factor $\alpha_{i}=\frac{1}{M}$ for each client. We show that the upper limit of the average information-theoretic distributed learning gap $\frac{1}{M}\big{|}\mathcal{L}_{Z^{\mathcal{I}_{p}}}(\hat{h})-\mathcal{L}_{{Z_{\mathcal{I}_{p}}}}(\hat{h})\big{|}$ is related to the entropy rate of the stochastic process $\{Z_{i}\}_{i\in\mathcal{I}_{p}}$.
@@
$=\big{|}\mathbb{E}_{Z^{\mathcal{I}_{p}}}\frac{\sum_{i\in\mathcal{I}_{p}}\ell(\hat{h}_{t},Z_{i})\log\frac{1}{P(Z^{\mathcal{I}_{p}})}}{M}-\frac{1}{K}\sum_{i\in\mathcal{I}_{t}}\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\ell(\hat{h}_{t},s_{i}^{j})\big{|},$
@@
$|\mathcal{L}_{Z^{\mathcal{I}_{p}}}(\hat{h}_{t})-\mathcal{L}_{S_{t}}(\hat{h}_{t})|\leq$ (12)
$b(3-\frac{2K}{M}-\frac{1}{K})H(Z^{\mathcal{I}_{p}})+\frac{b}{K}\sum_{i\in\mathcal{I}_{p}\setminus\mathcal{I}_{t}}H(P_{Z_{i}},P_{Z_{j}})+\mathcal{E}_{t}$
@@ IV-A Maximum Entropy Aggregation
Privacy Computations and Communication Costs: Privacy computation techniques can be integrated into the proposed methods: clients upload the empirical entropy of local data sources via homomorphic encryption or secure multi-party computation with low computation overhead since the empirical entropy is a scalar value. We assume that data sources of participating clients are stationary during the whole FL process and hence clients only need to upload the empirical entropy to the server before FL starts only for one time, which can reduce the communication cost.
@@ IV-B Gradient similarity-based client selection
where $\nabla F_{i}(\mathbf{w})$ denotes the practical local gradient calculated by client $i$ without considering the self information of samples, i.e., $\nabla F_{i}(\mathbf{w})=\nabla\ell(\mathbf{w},Z_{i})$, where $\mathbf{w}$ denotes the global model and $Z_{i}$ denotes the $i$-th data source. $\sigma_{\mathcal{K}}^{2}$ depends on the dissimilarity of data distributions $\{P_{Z_{k}}\}_{k\in\mathcal{K}}$. In other words, a higher level of distribution dissimilarity of $\{P_{Z_{k}}\}_{k\in\mathcal{K}}$ will incur a higher level of $\sigma_{\mathcal{K}}^{2}$ [33, 34]. Hence, $\sigma^{2}_{\mathcal{K}}$ is related to the term $\sup_{i,j\in\mathcal{K}}KL(P_{Z_{i}}\|P_{Z_{j}})$.
@@ IV-B1 Minimax gradient similarity-based client selection
4             $S_{i,j}\leftarrow S(\mathbf{g}_{t}^{i},\mathbf{g}_{t}^{j})$;
5             $\mathbb{S}_{i}\leftarrow\mathbb{S}_{i}\cup\{S_{i,j}\}$
@@
7      $\mathbb{S}_{max}\leftarrow\mathbb{S}_{max}\cup\{S_{i}^{*}=\mathop{\max}\limits_{S_{i,j}}\mathbb{S}_{i}\}$
@@
9while $|\tilde{\mathcal{I}_{t}}|\leq|\mathcal{I}_{t}|$ do
@@ IV-B2 Convex hull construction-based client selection
We outline the general procedure of two proposed client selection methods in Algorithm 2.
@@ V Experiment
In this section, we evaluate the proposed methods on three common datasets in FL, in order to verify our theoretical results. More details of experimental settings and results are provided in the appendix.
@@
Experiment setting: We consider three datasets commonly used in FL: i) image classification on EMNIST-10 and CIFAR-10 with a CNN model, ii) next character prediction on Shakespeare with a RNN model. For the image classification task, we split the dataset into different clients using the Dirichlet distribution spitting method, and we compare the proposed weighting aggregation method and client selection methods with baselines provided below. For the Shakespeare task, each speaking role in each play is set as a local dataset [38], and we only compare the proposed client selection methods with baselines on this dataset. We split all datasets into $100$ clients, and randomly select $40$ clients among them as participating clients. The remaining $60$ clients are considered non-participating clients. We evaluate the global model’s performance on two metrics: In Distribution (ID) performance and Out-of-Distribution (OOD) performance. The ID performance evaluates the global model on the local test set of selected clients, while the OOD performance evaluates the global model on a standard test set with a distribution equivalent to the total dataset of $100$ clients. All the selected clients perform $5$ local epochs before sending their updates. We use a batch size of $128$ and tune the local learning rate $\eta$ over a $\{0.1,0.01,0.001\}$ grid in all experiments.

### v5

@@ I Introduction
Federated Learning (FL) offers a collaborative paradigm to train a shared global model across distributed clients, ensuring data privacy by eliminating the need for direct data transfers [1, 2, 3]. Therefore, FL provides a secure architecture to bolster confidentiality and effectively manage diverse sensitive data [4, 5], ranging from financial to healthcare information. Consequently, FL applications span various domains, encompassing finance [6], healthcare [4], recommendation systems [7], Internet-of-Things (IoT) [8], and more. However, the inherent heterogeneity among clients—often due to their distinct operational environments—generates Non-Independent and Identically Distributed (Non-IID) scenario [9, 10, 11]. This distinctness complicates the assessment of FL’s generalization capabilities, setting it apart from traditional centralized learning [12, 13, 14].
@@
Motivation. Compared with traditional FL deploying a global model for actively participating clients and assuming an alignment between training and testing data distributions, our investigation spotlights the global model’s out-of-distribution generalization capabilities when trained by participating clients. In essence, we seek to cultivate a model via FL across active clients that reliably serves even the passive ones. Specifically, a novel contribution of our paper is the introduction of the self-information weighted expected risk – a metric to gauge model generalization. Our hypothesis is grounded in the conception that a model exhibiting proficiency with low-probability examples from training distributions might demonstrate adaptability to unfamiliar testing distributions. Such examples are anticipated to hold a greater significance in these unseen datasets. Operationalizing this concept within FL, our framework leverages the self-information of examples to craft a generalization boundary. Delving into the resulting information entropy-aware and distribution discrepancy-aware generalization disparities, a revelation emerges: data sources tend to manifest informational redundancy. This implies that certain data sources, characterized by diminished informational weight, can be seamlessly supplanted by others. Essentially, only a fraction of the client base holds substantial sway over FL’s generalization. This redundancy is especially pronounced in the IoT scenario, where a myriad of edge devices—often operating in overlapping zones—partake in FL [21]. Consider drones, serving as FL clients, and amassing spatial data for model training. Drones constrained by limited flight spans might be rendered redundant due to their shared operational territories with other drones [22]. Informed by these insights, our paper further proposes strategies to enhance the generalization capabilities of FL, enabling trained models to better serve non-participating clients.
@@ II Related Work
Besides, compared with Federated Domain Generalization (FedDG) [14, 31, 32] dealing with the challenge of domain distribution shift under FL settings, our focus is solely on a specific collection of all the data sources in FL. Our objective is to train a global model from seen distributions of participating clients and generalize it to other unseen distributions of non-participating clients. Additionally, it is commonly assumed that each domain contains equal information and cannot represent each other in FedDG. In other words, FedDG does not consider the inherent distribution correlation or information redundancy across data sources. However, we consider the presence of information redundancy among data sources and only a subset of clients contributes to the generalization of FL. Our focus is thus to identify representative data sources. In terms of developed algorithms, most DG studies concentrate on learning an invariant representation across different domains [33, 34, 35]. Conversely, our focus is to design proper weighting aggregation and client selection methods to mitigate the generalization error in FL.
@@ III Theoretical Framework
Similar to previous studies [10, 12], we model each data source as a random variable with its corresponding distribution. In order to better evaluate the generalization performance of FL to ensure trained models perform well for non-participating clients, we focus on the information-theoretic generalization gap in FL. Furthermore, similar to [10, 13], we also utilize uniform convergence-based generalization analysis in this study. This technique, as discussed by [36], is algorithm-independent, making it applicable to a wide range of federated optimization algorithms [2, 37, 38]. To conserve space, we have provided proofs for all our theorems in the appendix.
@@ III-A Preliminaries
The local training set $S_{i}=\{s_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in\mathcal{I}_{p}$ stored on participating client $i$ is made of $n_{i}$ i.i.d. realizations from the local data source $Z_{i}\sim P_{Z_{i}}$. Referring to [12], the objective of federated generalization is to train a global model on $\{S_{i}\}_{i\in\mathcal{I}_{p}}$, such that all the possible clients in $\mathcal{I}$ will be provided satisfactory service by this global model trained by participating clients. Let $\mathcal{H}$ be a hypothesis class on $\mathcal{X}$. The loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ is a non-negative function and we assume that $\ell$ is bounded by $b$ and Lipschitz continuous. For simplicity, we denote $\ell(h(x),y)$ as $\ell(h,z)$ in the following.
@@
The rationale behind the formulation of the risk outlined in (1) can be attributed to the fact that target distributions are unknown in the training stage under the OOD setting. The proposed loss $\ell(h,z_{i})\log(\frac{1}{P_{Z_{i}}(z_{i})})$, is indicative of the requisite focus on outcomes $z_{i}\in\mathcal{Z}$ with lower probabilities in source distributions since they may have higher probabilities in unknown target distributions. Additionally, it is reasonable to assume that unknown data sources have maximum entropy distributions, signifying greater uncertainty, exemplified by a uniform distribution of discrete labels. For example, IoT devices may only collect data in specific areas, while the global model trained by these devices is expected to provide spatial-related services for all devices across the entire area [21]. Hence, the self-information of each outcome is indispensable for measuring the expected risk in such a situation. Moreover, we know that applying FL in the healthcare field [39] should focus on silos within rare disease cases, which underscores the significance of rare yet informative samples, thereby supporting our viewpoint.
@@ III-B Federated Generalization
Theorem 1 asserts that increasing the weighted sum $\sum_{i\in\mathcal{I}_{p}}\alpha_{i}H(Z_{i})$ will reduce this generalization bound. This suggests that models trained by FL will exhibit enhanced performance on unknown data sources when a greater weighting factor is assigned to the data source with richer information. In addition, model complexity $VC(\mathcal{G})$ and sample complexity $\sum_{i=1}^{|\mathcal{I}_{p}|}n_{i}$ also affect the generalization capacity of FL. Notice that alternative metrics, like Rademacher complexity and covering number, can be employed to refine VC-dimension-based generalization bounds [10, 40], but delving into these is beyond the scope of this paper. Our emphasis lies in devising algorithms by leveraging insights from the proposed information-theoretic generalization framework.
We then consider assigning the identical weighting factor $\alpha_{i}=\frac{1}{M}$ for each client. To explore the impact on each individual client resulting from the distributed learning performed on decentralized data sources, we examine the average information-theoretic distributed learning gap defined as $\frac{1}{M}\big{|}\mathcal{L}_{Z^{\mathcal{I}_{p}}}(\hat{h})-\mathcal{L}_{{Z_{\mathcal{I}_{p}}}}(\hat{h})\big{|}$ and show that its upper bound is related to the entropy rate of the stochastic process $\{Z_{i}\}_{i\in\mathcal{I}_{p}}$.
@@
$=\Big{|}\mathbb{E}_{Z^{\mathcal{I}_{p}}}\frac{\sum_{i\in\mathcal{I}_{p}}\ell(\hat{h}_{t},Z_{i})\log\frac{1}{P(Z^{\mathcal{I}_{p}})}}{M}-\sum_{i\in\mathcal{I}_{t}}\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\frac{\ell(\hat{h}_{t},s_{i}^{j})}{K}\Big{|},$
@@
$\leq b(3-\frac{2K}{M}-\frac{1}{K})H(Z^{\mathcal{I}_{p}})+\frac{b}{K}\sum_{i\in\mathcal{I}_{p}\setminus\mathcal{I}_{t}}H(P_{Z_{i}},P_{Z_{j}})+\mathcal{E}_{t}$
@@ IV-A Maximum Entropy Aggregation
In the following, we will discuss the effect of privacy computation, noisy data and computation-communication costs on our proposed method.
Privacy Computation: Privacy computation techniques can be integrated into the proposed methods: clients upload the empirical entropy of local data sources via homomorphic encryption or secure multi-party computation with low computation overhead since the empirical entropy is a scalar value.
Noisy Data Challenges: If local models are trained on severely noisy data, we can regard such challenge as a form of Byzantine attack [42] and address this issue using corresponding solutions. We can also identify outliers with extremely large entropy and expel them from FL. Additionally, replacing entropy with a relative metric of calculating the similarity between the local label distribution and uniform distribution can mitigate the effect from noisy data. Besides, the methods of Noisy Clients Detection and Noisy Robust Aggregation proposed in [43] can be integrated with our own approaches to alleviate the influence of noisy data.
Computation and Communication Costs: In this study, we assume that data sources of participating clients are stationary during the whole FL process. Hence, clients only need to upload empirical entropy to the server for one time, which can reduce the communication cost.
@@ IV-B Gradient similarity-based client selection
where $\nabla F_{i}(\mathbf{w})$ denotes the practical local gradient calculated by client $i$ without considering the self information of samples, i.e., $\nabla F_{i}(\mathbf{w})=\nabla\ell(\mathbf{w},Z_{i})$, where $\mathbf{w}$ denotes the global model and $Z_{i}$ denotes the $i$-th data source. $\sigma_{\mathcal{K}}^{2}$ depends on the dissimilarity of data distributions $\{P_{Z_{k}}\}_{k\in\mathcal{K}}$. In other words, a higher level of distribution similarity of $\{P_{Z_{k}}\}_{k\in\mathcal{K}}$ will incur a lower level of $\sigma_{\mathcal{K}}^{2}$ [44, 37]. This is because that $\|\nabla\mathcal{L}_{P_{z_{i}}}(z,w)-\nabla\mathcal{L}_{P_{z_{j}}}(z,w)\|^{2}=\|\int_{\mathcal{Z}}\nabla\ell(w,z)(dP_{z_{i}}-dP_{z_{j}})\|^{2}\leq c^{\prime}KL(P_{z_{i}}\|P_{z_{j}})$, the last inequality holds follows from Lipschitz continuous loss and Pinsker’s inequality. Hence, $\sigma^{2}_{\mathcal{K}}$ is related to the term $\sup_{i,j\in\mathcal{K}}KL(P_{Z_{i}}\|P_{Z_{j}})$.
@@
Notice that the server can maintain a dynamic table containing the client IDs, their respective entropy values, and the latest local gradients. This table empowers the server to effectively handle dynamic clients and execute proposed algorithms. Hence, the proposed method is robust to client churn, such as clients frequently joining, leaving, and rejoining the system.
@@ IV-B1 Minimax gradient similarity-based client selection
4             $S_{i,j}\leftarrow S(\mathbf{g}_{t}^{i},\mathbf{g}_{t}^{j})$ // Similarity computation;
5             $\mathbb{S}_{i}\leftarrow\mathbb{S}_{i}\cup\{S_{i,j}\}$ // Similarity set construction
@@
7      $\mathbb{S}_{max}\leftarrow\mathbb{S}_{max}\cup\{S_{i}^{*}=\mathop{\max}\limits_{S_{i,j}}\mathbb{S}_{i}\}$ ;
8      // Max similarity set construction
@@
10while $|\tilde{\mathcal{I}_{t}}|\leq|\mathcal{I}_{t}|$ do
@@ IV-B2 Convex hull construction-based client selection
We outline the general procedure of two proposed client selection methods in Algorithm 2. Then we provide some insights of reducing the communication and computation cost of these client selection methods.
Communication and Computation cost: To further reduce the communication and computation cost of proposed methods, we can employ the event-triggered communication techniques [48, 49] widely used in distributed optimization to diminish the overhead caused by frequent communication of local gradients. Moreover, based on event-triggered communication, the server can utilize historical gradient similarities stored in another table to avoid redundant computations.
@@ V Experiment
In this section, we first evaluate the proposed methods on three common datasets in FL, in order to verify our theoretical results. Then we compare the proposed methods with more baselines on CIFAR-100 dataset to further validate the generalization performance of our methods.
Experiment setting: We fisrt consider three datasets commonly used in FL: i) image classification on EMNIST-10 and CIFAR-10 with a CNN model, ii) next character prediction on Shakespeare with a RNN model. For the image classification task, we split the dataset into different clients using the Dirichlet distribution spitting method, and we compare the proposed weighting aggregation method and client selection methods with baselines provided below. For the Shakespeare task, each speaking role in each play is set as a local dataset [50], and we only compare the proposed client selection methods with baselines on this dataset. We split all datasets into $100$ clients, and randomly select $40$ clients among them as participating clients. The remaining $60$ clients are considered non-participating clients. We evaluate the global model’s performance on two metrics: In Distribution (ID) performance and Out-of-Distribution (OOD) performance. The ID performance evaluates the global model on the local test set of selected clients, while the OOD performance evaluates the global model on a standard test set with a distribution equivalent to the total dataset of $100$ clients. All the selected clients perform $5$ local epochs before sending their updates. We use a batch size of $128$ and tune the local learning rate $\eta$ over a $\{0.1,0.01,0.001\}$ grid in all experiments.
@@
Additional Experiments on CIFAR-100 with ResNet-18: To further assess the generalization performance of the proposed methods, we proceed with additional experiments, comparing our approaches with recent baselines using the CIFAR-100 dataset [52] and a ResNet-18 model [53]. Initially, we partition the dataset into 100 clients using a Dirichlet distribution split with a parameter $\alpha$ set to 0.03. We then randomly choose 40 clients from this set to participate in FL. Each selected client conducts local training with a batch size of 16, runs for 3 local epochs, and uses a learning rate of 0.01. Our focus remains exclusively on the OOD performance of the global model.
We compare the proposed empirical entropy-based weighting method with two baselines, namely FedALRC [40] and FedGAMMA [38]. Additionally, we compare two presented client selection methods with the active client selection proposed in [54]. The experimental results are depicted in Fig. 7.
(a) Client selection method
(b) Weighting method
Figure 7: The convergence analysis on OOD test accuracy of empirical entropy weighting method and two proposed client selection methods compared with baselines on CIFAR-100.
In Fig. 7 (a), we employ a third-order moving average to present the results. We observe that the proposed convex hull construction-based client selection method performs worse than active selection in the later stage of FL. However, the proposed selection method based on minimax gradient similarity outperforms this baseline. This difference in performance can be attributed to the sensitivity of building convex hulls for complex models such as ResNes models, while using cosine similarity to measure gradient similarity demonstrates increased robustness in this situation. Fig. 7 (b) illustrates that the proposed empirical entropy-based weighting method outperforms other baseline methods. This superiority is due to the fact that these baseline methods only consider enhancing I.I.D. generalization performance and overlook potential non-participating clients. In contrast, our presented empirical entropy-based weighting method can improve the generalization capacity of models by identifying representative data sources.